3.A bowl of raffle tickets contains the numbers 1 through 80. What is the probability that a randomly selected ticket will be a number that is not a multiple of 9?
Answer by greenestamps(13195)



Number of cards: 80

Number of cards numbered with numbers that ARE a multiple of 9: 8 (9, 18, 27, ..., 63, 72)

Number of cards that are numbered with numbers that are NOT a multiple of 9: 80-8 = 72

ANSWER: P(card is not a multiple of 9): 72/80 = 9/10 = 0.90 = 90%.